OpenStudy (anonymous):

I have tha answer can someone please just check me... Find the derivative of p(x)=cos(ln(3x-1))

jimthompson5910 (jim_thompson5910):

you got what?

OpenStudy (anonymous):

I got \[\frac{ 3\sin(\ln(3x-1)) }{ 3x-1 }\]

jimthompson5910 (jim_thompson5910):

again you lost a negative sign somewhere, the derivative of cosine is negative sine

jimthompson5910 (jim_thompson5910):

p(x) = cos(ln(3x-1)) p ' (x) = d/dx[ cos(ln(3x-1)) ] p ' (x) = -sin(ln(3x-1))*d/dx[ln(3x-1)] p ' (x) = -sin(ln(3x-1))/(3x-1)*d/dx[3x-1] p ' (x) = -sin(ln(3x-1))/(3x-1)*3 p ' (x) = -3sin(ln(3x-1))/(3x-1)

jimthompson5910 (jim_thompson5910):

other than that, everything else is correct

OpenStudy (anonymous):

ok so its just a negative on the top then

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np